I need to know how to subtract fractions of a number.
My problem is 15 - 2/3.
Thanks

Answer by checkley77(12569)

15 - 2/3.
15-2/3=14 1/3
or
(15*3-2)/3=(45-2)/3=43/3 or 14 1/3
or
15=45/3 thus:
45/3
-2/3
-----------
43/3=14 2/3
-------------------------------------------------------------
Pick the method that is the easiest for you to understand.